Come-Clean Corporation produces a variety of cleaning compounds and solutions for both industrial and household use. While moRequired: 1. How much incremental revenue does the company earn per jar of polish by further processing Grit 337 rather thanCORRECT the WRONG answers
Answer #1

Solution 1:

Revenue per Jar of silver polish = $5 per jar

Revenue per 1/4th pound of Grit 337 = $7.20 / 4 = $1.80

Incremental revenue per jar for further processing = $5 - $1.80 = $3.20 per jar

Solution 2:

Computation of Additional contribution Margin- If Grit 377 sold after further processing
Particulars Grit 377 per 1/4 Pound Silver Polish per jar Net Increase/ (decrease)
Selling price per unit $1.80 $5.00 $3.20
Variable Costs:
Making cost of Grit 337 $0.30 $0.30 $0.00
Other ingredients $0.00 $0.55 $0.55
Direct labour $0.00 $1.44 $1.44
Variable Manufacturing Overhead $0.00 $0.36 $0.36
Variable Selling Cost $0.00 $0.25 $0.25
Total variable costs $0.30 $2.90 $2.60
Contribution margin $1.50 $2.10 $0.60

Hence incremental contribution per jar for further processing = $0.60 per jar

Solution 3:

Minimum number of jars of silver polish that must be sold each month = Avoidabe fixed cost / incremental contribution margin per jar

= ($3400+$2600) / $0.60 = 10,000 jars

Solution 4:

If company sells 8400 jars of polish then incremental contribution margin of further processing = 8400 * $0.60 = $5,040

Avoidable fixed cost = $3400 + 2600 = $6,000

Financial advantage (disadvantage) of further processing = $5,040 - $6,000 = ($960)

Solution 5:

If company sells 11200 jars of polish then incremental contribution margin of further processing = 11200 * $0.60 = $6,720

Avoidable fixed cost = $6,000

Financial advantage (disadvantage) of further processing = $6,720 - $6,000 = $720